# references.bib

### v1

@@ I INTRODUCTION
According to the World Health Organization, there are 253 million people estimated to be visually impaired and 36 million people totally blind in the world [Bourne2017]. Blind people face difficulties navigating the outdoors, engaging in social activities, browsing the internet, or even everyday activities, like currency denomination or house chores. So, visual impairment poses a significant challenge to the perception and interpretation of environmental stimuli, showcasing the need for auxiliary technological aids. Assistive technology for the visually impaired has seen great advancements in recent years, including electronic travel aids, like Smart Cane [SmartCane], Bbeep [Kayukawa2019] or Shoe [shoe], Orientation Aids like Smart Walker [SmartWalker] and even Position Locator Devices [Manjari2020]. However, current assistive devices offer limited functionality, failing to
Figure 1: AIris architecture overview: User is wearing the device including earphones, camera and Raspberry Pi. The user asks a question, which Raspberry Pi converts to text and using NLP it selects which module should be evoked. The camera input is transmitted to the server, where the ML model inference happens. The results are sent back to the Raspberry Pi and the final answer is formatted converted from speech to text and communicated through the earphones.
provide comprehensive spatial awareness and independence [Manjari2020].
@@
In this study, we develop AIris, an assistive system for the visually impaired with AI-enhanced functionalities. AIris uses state-of-the-art machine learning models to facilitate environmental understanding through auditory feedback. We design and build a prototype of this assistive device, using a pair of 3D-printed glasses equipped with a camera slot, earphones, and a microcomputer. We develop a fully operational prototype system that can be used in varying real-life scenarios and daily activities. Using computer vision techniques and advanced pre-trained neural networks, it analyzes camera data to provide comprehensive feedback about face detection, text reading, object recognition, money counting, and many more. AIris functionalities aim to makes the environment more accessible and give visually impaired people greater independence in their everyday activities.
@@ II-A System Overview
The hardware infrastructure of AIris is designed to fulfil the needs of visually impaired users. We use a distributed architecture, capturing local information and transmitting to a central server for computationally-intensive processing (Fig. 1). The device’s primary sensory interface is a camera, centrally positioned on the eyewear, capturing high-resolution visual data from the user’s perspective. This visual information is initially processed by a Raspberry Pi microcomputer, serving as a local computational node.
The microcomputer then securely transmits the data to a centralized server, where intensive image processing and analysis are conducted. The server, leveraging superior computational resources, applies advanced machine learning models for tasks such as object detection, facial recognition, and text extraction. Upon completing the analysis, the server dispatches the processed results back to the Raspberry Pi.
On receipt of the server’s data, the Raspberry Pi undertakes the final assembly of the response, using a natural language processing (NLP) engine. The NLP engine is responsible for translating analytical data into intelligible auditory feedback, which is then communicated to the user via the earphones. This processing approach —local data acquisition and remote analysis— ensures that AIris provides rapid and accurate environmental descriptions, enhancing the user’s spatial awareness with minimal latency.
@@ II-B1 Face Recognition Module
This module detects and recognizes human faces within the camera’s visual field (Fig. 2)[FaceRecognition]. It checks recognized faces against a user-defined dynamic database and tells the user who the person is if known, otherwise, the users can choose to add the new faces to the system. This is essential for personalizing the user experience, as it provides the ability to remember and identify acquaintances, family members, and friends.
@@ II-B3 Text Reading Module
The text reading module is a critical component of the AIris system, serving as a bridge between the visually impaired and the vast world of written communication. Utilizing state-of-the-art optical character recognition (OCR) technology, this module scans the visual field for textual content, interprets it, and converts it into spoken words[Tesseract]. This module enables the visually impaired to access the vast textual information in labels, instructions, signs and documents.
@@ II-B4 Object Recognition Module
The object recognition module harnesses sophisticated machine learning algorithms to identify and classify objects within the camera’s field of view (Fig. 2) [Yolo]. The module is particularly significant as it allows users to interact with their environment in a more meaningful way, recognizing items that are used in daily life, from personal belongings to potential hazards.
@@ II-B5 Counting Money
This module is designed to specifically address the challenges visually impaired individuals face when handling currency (Fig. 3). This module employs a combination of image processing and pattern recognition algorithms to differentiate and count banknotes and coins. Its importance is underscored by the critical need for financial independence and the confidence to conduct transactions without assistance.
@@ II-B6 Note Taking Module
This module utilizes advanced speech recognition technology to convert spoken words into digital text, which can then be saved and accessed later as needed Fig. 3). The significance of this module emerges from its facilitation of personal organization and memory assistance, which are essential for maintaining independence.
@@ II-C2 Voice Commands Interface
Upon system activation, the user is invited to communicate with AIris using natural language. This interaction is facilitated by an advanced natural language processing (NLP) engine [GoogleSpeech]. Users issue commands or queries in natural language, such as requesting a description of their environment, identifying an object, or asking for the time. The NLP engine analyzes the speech for intent and relevant information, so it finally maps the spoken phrases to predefined operations that the AIris system can execute.
@@
Figure 4: AIris Design Process: This figure shows from left to right the evolution of the hardware for AIris. At first, we sketched out the basic idea and how everything would interact. Then we moved on to designing the eyewear in a Computer Aided Design (CAD) Program, focusing on keeping it light-weight, simple and user-friendly. Finally, we 3D printed our eyewear and assembled the whole prototype.
II-C3 Custom Eye Wear Design
The eyewear is designed meticulously in order to avoid obstructing the field of view of low-vision individuals. The objective was to develop a lightweight, comfortable frame capable of supporting a camera mount without compromising the user’s experience. The camera mount was engineered to position the camera at the center of the glasses, mirroring the natural human line of sight. This positioning is critical for accurate data capture corresponding to the wearer’s perspective. Various prototypes were printed to test the balance, fit, and comfort of the eyewear, with adjustments made to the CAD models in response to user feedback. The design process is illustrated through the iterations in Fig. 4.
@@ II-C4 Hardware Integration
In the prototypical stage of AIris, the thoughtful integration of hardware components has been a critical focus to ensure functionality while maintaining user comfort. The camera, a compact yet powerful device, is strategically mounted at the front center of the eyewear. It is connected via a flexible, lightweight cable to the Raspberry Pi microcomputer, which is conveniently housed within the user’s pocket. This arrangement keeps the eyewear as light as possible, reducing the strain on the user’s nose and ears during prolonged use.
The Raspberry Pi microcomputer, selected for its compact design and powerful processing capabilities, processes visual data at high speeds with minimal power consumption.
The Bluetooth earphones, chosen for their lightweight and wireless nature, are paired seamlessly with the system, whereas the integration process involved optimizing the Bluetooth connectivity to ensure a stable and consistent link between the earphones and the Raspberry Pi.
@@ II-C5 Local Software & Raspberry Pi
The local software running on the Raspberry Pi within the AIris system is chosen with respect to efficacy and performance [RaspberryPi]. The first step of the user interaction is the speech-to-text conversion, facilitated by Google’s Speech Recognition Engine [GoogleSpeech]. This service translates spoken language into textual commands with very high accuracy, even in noisy environments, while also maintaining minimal latency.
@@
When textual content needs to be deciphered from the environment, the system leverages pytesseract, a Python wrapper for Google’s Tesseract-OCR Engine [Tesseract]. Pytesseract is adept at recognizing text within a wide variety of lighting conditions and font styles, converting visual text into machine-encoded text.
The counting money module combines computer vision filters with optical character recognition (OCR), drawing inspiration from [money] to facilitate the identification and counting of currency. Upon receiving visual data of money, the server first applies specialized filters to enhance the features of the banknotes or coins, making them more distinguishable. Following this enhancement, the pytesseract OCR engine is employed to read the denominations. The module is adept at handling various currencies, orientations, and lighting conditions, providing users with an accurate count and denomination of their money.
The note-taking module is a significant feature that enhances the user’s ability to record and organize information. This module operates by converting spoken words into text using Google’s Speech Recognition engine [GoogleSpeech]. Once the speech is transcribed, the system categorizes the text based on user-defined categories such as reminders, notes, or lists. The categorized notes are then saved into files which can be easily retrieved and reviewed by the user.
To achieve efficient and accurate barcode scanning, this module leverages Pyzbar, a python library known for its proficiency in decoding barcodes and QR codes[PyZBar]. When the user points the AIris camera at a barcode, Pyzbar processes the image to detect and decode the barcode. Given the varying qualities of barcodes and the diverse conditions under which they might be scanned, the server employs advanced image processing techniques to enhance the barcode’s features. These preprocessing steps include adjusting brightness and contrast, sharpening, and scaling to ensure that Pyzbar can accurately decode the information even in less-than-ideal circumstances.
@@ II-C6 Server Software
For the scene description module, we deploy MAX (Model Asset eXchange), an Image Caption Generator [Max]. This model generates captions from a fixed vocabulary that describe the contents of images in the COCO Dataset [COCO]. The model consists of an encoder model - a deep convolutional net using the Inception-v3 architecture trained on ImageNet-2012 data - and a decoder model - a network that is trained on the encoding from the encoder model [Inception], [imagenet]. The input to the model is an image, and the output is a sentence describing the image content.
@@ III RESULTS & DISCUSSION
  • ’read’ all digital & printed text through audio output, making every menu, sign or document accessible to them
@@
  • verbally record notes, aiding in personal organization
@@
Each of those features transforms visual data into auditory information, helping visually impaired users interact with their environment. AIris’ feedback is immediate and can improve the users’ independence, helping with many daily tasks. Examples of the user-system interaction can be seen in Figure 2 and Figure 3, as a demonstration of the fully integrated system in use.
To measure AIris performance, we can note the machine learning models’ accuracy in popular datasets, like:
Model/Library Performance
Face Recognition - face_recognition (ResNet-34)  99.38% accuracy on LFW dataset [LFWTech]
Object Detection - YOLO (You Only Look Once)  63.4% mean Average Precision on COCO dataset [COCO]
Speech to Text - Google’s Speech Recognition  95% accuracy for English
Scene Description - MAX Image Caption Generator  27% BLEU [bleu] score on Flickr8k dataset [flickr8k]
@@
AIris, in its current prototype form, has demonstrated significant potential in aiding visually impaired users. AIris modules have been tested in an unofficial trial. This trial was instrumental in evaluating the practical applicability and user experience of AIris in real-world scenarios.
The trial involved three visually impaired individuals, using AIris in a controlled environment. The participants were asked to engage in various tasks such as paying for something, identifying objects, reading text, and recognizing faces. Their feedback was collected to measure the effectiveness of AIris and identify areas for improvement.
@@
AIris doesn’t require a lot of resources since heavy machine learning models are running on the cloud. The system is flexible, updateable, and upgradeable since each compartment (module) is easily replaced with a new one. By situating the microcomputer in the user’s pocket, we effectively distribute the system’s weight, making the eyewear feel as close to ordinary glasses as possible. AIris is designed to be module-based, so it can incorporate new features and functionalities. We can add new state-of-the-art machine learning models and expand AIris functionalities with new modules. Finally, both hardware and software can be personalized, from using different cameras and designing different styles of glasses to creating custom personalized features according to each user’s needs.

### v2

@@ I INTRODUCTION
According to the World Health Organization, there are 253 million people estimated to be visually impaired and 36 million people totally blind in the world [Bourne2017]. Blind people face difficulties navigating the outdoors, engaging in social activities, browsing the internet, or even everyday activities, like currency denomination or house chores. So, visual impairment poses a significant challenge to the perception and interpretation of environmental stimuli, showcasing the need for auxiliary technological aids. Assistive technology for the visually impaired has seen great advancements in recent years, including electronic travel aids, like Smart Cane [SmartCane], Bbeep [Kayukawa2019] or Shoe [shoe], Orientation Aids like Smart
Figure 1: AIris architecture overview: User is wearing the device, including earphones, camera and Raspberry Pi. The user asks a question, which Raspberry Pi converts to text and using Natural Language Processing it selects which module should be evoked. The camera input is transmitted to the server, where the ML model inference happens. The results are then sent back to the Raspberry Pi, where they are formatted into a sentence, converted from text to speech, and finally communicated to the user through the earphones.
Walker [SmartWalker] and even Position Locator Devices [Manjari2020]. However, current assistive devices offer limited functionality, failing to provide comprehensive spatial awareness and independence [Manjari2020].
@@
In this study, we develop AIris, an assistive system for the visually impaired with AI-enhanced functionalities. AIris uses state-of-the-art machine learning models to facilitate environmental understanding through auditory feedback. We design and build a prototype of the assistive device, by designing a custom pair of 3D-printed glasses equipped with a camera slot, earphones, and a microcomputer. We develop a fully operational prototype system that can be used in varying real-life scenarios and daily activities. Using computer vision techniques and advanced pre-trained neural networks, it analyzes camera data to provide comprehensive feedback about face detection, text reading, object recognition, money counting, and many more. AIris functionalities aim to make the environment more accessible and give visually impaired people greater independence in their everyday activities.
@@ II-A System Overview
The hardware infrastructure of AIris is designed to fulfill the needs of visually impaired users. We use a distributed architecture, capturing local information and transmitting to a central server for computationally-intensive processing (Fig. 1). The device’s primary sensor is a camera, centrally positioned on the eyewear, capturing high-resolution visual data from the user’s perspective. This visual information is initially processed by a Raspberry Pi microcomputer, serving as a local computational node.
The microcomputer then securely transmits the data to a centralized server, where intensive image processing and analysis are conducted. The server, leveraging superior computational resources, applies advanced machine learning models for tasks such as object detection, scene description, and text extraction. Upon completing the analysis, the server dispatches the processed results back to the Raspberry Pi.
On receipt of the server’s data, the Raspberry Pi undertakes the final assembly of the response using a natural language processing (NLP) engine. The NLP engine is responsible for translating analytical data into intelligible auditory feedback, which is then communicated to the user via the earphones. This processing approach —local data acquisition and remote analysis— ensures that AIris provides rapid and accurate environmental descriptions, enhancing the user’s spatial awareness with minimal latency.
@@ II-B1 Face Recognition Module
This module detects and recognizes human faces within the camera’s visual field (Fig. 2) [FaceRecognition]. It checks recognized faces against a user-defined dynamic database and tells the user who the person is if known, otherwise, the users can choose to add the new faces to the system. This is essential for personalizing the user experience, as it provides the ability to remember and identify acquaintances, family members, and friends. To safeguard user privacy, we exclusively store facial landmark embeddings, eliminating the need to retain the full image. Due to the low computational requirements and to ensure low latency, this module is implemented on the Raspberry Pi, running locally.
@@ II-B3 Text Reading Module
The text reading module is a critical component of the AIris system, serving as a bridge between the visually impaired and the vast world of written communication. Utilizing state-of-the-art optical character recognition (OCR) technology, based on Google’s Tesseract, this module scans the visual field for textual content, interprets it, and converts it into spoken words [Tesseract]. This module enables the visually impaired to access the vast textual information in labels, instructions, signs and documents.
@@ II-B4 Object Recognition Module
The object recognition module harnesses a pre-trained version of Yolo [Yolo] to identify and classify objects within the camera’s field of view (Fig. 2). The module is particularly significant as it allows users to interact with their environment in a more meaningful way, recognizing items that are used in daily life, from personal belongings to potential hazards.
@@ II-B5 Counting Money
This module employs a combination of custom image processing and pattern recognition algorithms to differentiate and count banknotes and coins (Fig. 3). Its importance is underscored by the critical need for financial independence and confidence to conduct transactions without assistance.
@@ II-B6 Note Taking Module
This module utilizes advanced speech recognition technology to transcribe spoken words, save them, and access them later as needed (Fig. 3). The significance of this module emerges from its facilitation of personal organization and memory assistance, which are essential for maintaining independence.
@@ II-C2 Voice Commands Interface
Upon system activation, the user is invited to communicate with AIris using natural language. This interaction is facilitated by an advanced NLP engine, Google Speech [GoogleSpeech]. Users issue commands or queries in natural language, such as requesting a description of their environment, identifying an object, or reading visible text. The NLP engine analyzes the speech, mapping the spoken phrases, using defined keywords, to predefined operations that the AIris system can execute.
@@
Figure 4: AIris Design Process: This figure shows from left to right the evolution of the hardware for AIris. At first, we sketched out the basic idea and how the components would interact. Then we moved on to designing the eyewear in a Computer Aided Design (CAD) Program, focusing on keeping it light-weight, simple and user-friendly. Finally, we 3D printed our eyewear and assembled the whole prototype.
II-C3 Custom 3D Printed Eye Wear Design
The eyewear is designed meticulously in order to avoid obstructing the field of view of low-vision individuals. The objective was to develop a lightweight, comfortable frame capable of supporting a camera mount. The camera mount was engineered to position the camera at the center of the glasses, mirroring the natural human line of sight. Various prototypes were 3D printed to test the balance, fit, and comfort of the eyewear, with adjustments made to the CAD models in response to user feedback. The design process is illustrated through the iterations in Fig. 4.
@@ II-C4 Hardware Integration
In the prototypical stage of AIris, the thoughtful integration of hardware components has been a critical focus to ensure functionality while maintaining user comfort. The camera is strategically mounted at the front center of the eyewear. It is connected via a flexible, lightweight cable to the Raspberry Pi microcomputer, which is conveniently placed within the user’s pocket. This arrangement keeps the eyewear as light as possible, reducing the strain on the user’s nose and ears during prolonged use.
The Raspberry Pi 4B microcomputer, selected for its compact design and powerful processing capabilities, processes visual data at high speeds with minimal power consumption. To ensure extended mobility, the microcomputer utilizes a slim 20,000mAh power bank, providing over 6 hours of uninterrupted use.
The Bluetooth earphones are paired seamlessly with the system, while also enabling ambient mode, enhancing the user’s surroundings’ sounds so as to not reduce their hearing.
@@ II-C5 Local Software & Raspberry Pi
The local software running on the Raspberry Pi within the AIris system is chosen with respect to efficacy and performance [RaspberryPi]. The first step of the user interaction is the speech-to-text conversion, facilitated by Google’s Speech Recognition Engine [GoogleSpeech]. This service translates spoken language into textual commands with very high accuracy, even in noisy environments like a supermarket, or a cafeteria, while also maintaining minimal latency.
@@
When textual content needs to be extracted from the environment, the system leverages pytesseract, a Python wrapper for Google’s Tesseract-OCR Engine [Tesseract]. Pytesseract is adept at recognizing text within a wide variety of lighting conditions and font styles.
The counting money module utilizes computer vision filters and optical character recognition (OCR) for identification and counting of currency, drawing inspiration from the work detailed in [money]. Upon receiving the camera input, the server applies filters and thresholding to enhance the features of the banknotes or coins. Following this enhancement, the pytesseract OCR engine is employed to read the denominations. The module is adept at handling various currencies, orientations, and lighting conditions, providing users with an accurate count and denomination of their money.
The note-taking module is a significant feature that enhances the user’s ability to record and organize information. This module operates by converting spoken words into text using Google’s Speech Recognition engine [GoogleSpeech]. Once the speech is transcribed, the system categorizes the text based on user-defined categories such as reminders, notes, or lists. The categorized notes are then saved into text files which can be easily retrieved and reviewed by the user.
To achieve efficient and accurate barcode scanning, this module leverages Pyzbar, a python library known for decoding barcodes and QR codes[PyZBar]. When the user points the AIris camera at a barcode, Pyzbar processes the image to detect and decode the barcode. Given the varying qualities of barcodes and the diverse visual conditions, the server employs image processing techniques to enhance the barcode’s features. These preprocessing steps include adjusting brightness and contrast, sharpening, and scaling to ensure that Pyzbar can accurately decode the information even in less-than-ideal circumstances.
@@ II-C6 Server Software
For the scene description module, we deploy MAX (Model Asset eXchange), an Image Caption Generator [Max]. This model generates captions from a fixed vocabulary that describe the contents of images in the COCO Dataset [COCO]. The model consists of an encoder model - a deep convolutional net using the Inception-v3 architecture trained on ImageNet-2012 data - and a decoder model [Inception], [imagenet]. The input to the model is an image, and the output is a sentence describing the image content.
@@ III RESULTS & DISCUSSION
  • read digital & printed text through audio output, making every menu, sign or document accessible
@@
  • record verbal notes, aiding in personal organization
@@
Each of those features transforms visual data into auditory information, helping visually impaired users interact with their environment. AIris’ feedback is immediate and can improve the users’ independence, helping with many daily tasks. Examples of the user-system interaction can be seen in Figure 2 and Figure 3.
To measure AIris performance, we can note the machine learning models’ accuracy in popular datasets I. The high performance metrics and relatively low response times ensure a natural interaction with the aid.
Model/Library Performance Response Time
Face Recognition - face_recognition (ResNet-34)  99.38% accuracy on LFW dataset [LFWTech] 50ms
Object Detection - YOLO (You Only Look Once)  63.4% mean Average Precision on COCO dataset [COCO] 150ms
Speech to Text - Google’s Speech Recognition  95% accuracy for English 200ms
Scene Description - MAX Image Caption Generator  27% BLEU [bleu] score on Flickr8k dataset [flickr8k] 150ms
@@
AIris, in its current prototype form, has demonstrated significant potential in aiding visually impaired users. AIris modules have been tested in an unofficial trial to evaluate the practical applicability and user experience of AIris in real-world scenarios.
The trial involved three visually impaired individuals using AIris in a controlled environment. The participants were asked to engage in various tasks such as paying for something, identifying objects, reading text, and recognizing faces. Their feedback was collected to measure the effectiveness of AIris and identify areas for improvement.
@@
AIris doesn’t require a lot of resources since heavy machine learning models are running on the cloud. The system is flexible, updateable, and upgradeable. It can incorporate new features and functionalities, like new state-of-the-art machine learning models. Also, by situating the microcomputer in the user’s pocket, we effectively distribute the system’s weight, making the eyewear feel as close to ordinary glasses as possible. Finally, both hardware and software can be personalized, from using different cameras and designing different styles of glasses to creating custom personalized features according to each user’s needs.